OpenStudy (anonymous):

find the limit as h approaches 0: f(13+h) - f(13) divided by h if f(x) = ³√1695 - 8x^2

OpenStudy (anonymous):

\[ ³√1695 - 8x^2\] ?

OpenStudy (anonymous):

seems unlikely... missing something?

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ \sqrt[3]{1695 - 8x^2} -7 }{ h } \]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

same gimmick as with a square root, rationalize the numerator, but this time instead of using \[(a-b)(a+b)=a^2-b^2\] you have to use \[(a-b)(a^2+ab+b^2)=a^3-b^3\] so it is a pain in the arse

OpenStudy (anonymous):

the answer is \[\frac{ -208 }{ 147 }\] i want to figure what to do to get it.

OpenStudy (anonymous):

o i see

OpenStudy (anonymous):

have fun but you can do it use \(a=\sqrt[3]{1695-8x^2}\) and \(b=7\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \frac{ f(13 + h) - f(13) }{ h }\] this is the original question btw

OpenStudy (anonymous):

if f(x) is \[\sqrt[3]{1695 - 8x^2}\]

OpenStudy (anonymous):

what is \(f(13)\)?

OpenStudy (anonymous):

\[\sqrt[3]{1695 - 8x^2} \] with 13 plugged in, which equals 7

OpenStudy (anonymous):

ok then you can start with \[\frac{\sqrt[3]{1695-8(13+h)^2}-7}{h}\]\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

you can leave it in this form, or you can write \[\sqrt[3]{343-h^2-208h}-7\]

OpenStudy (anonymous):

who gave you this problem? this really sucks unless you are supposed to use a shortcut, namely recognize this as the derivative and evaluate

OpenStudy (anonymous):

university online homework

OpenStudy (anonymous):

mathXL

OpenStudy (anonymous):

do you know how to take a derivative? because then it it would be not so hard but if you do not, then there is a ton of work to be done

OpenStudy (anonymous):

i do know, im not sure if it would give me the answer im supposed to get

OpenStudy (anonymous):

ill try taking the derivative

OpenStudy (anonymous):

this is the derivative of \[\sqrt[3]{1695-8x^2}\] evaluated at \(x=13\)

OpenStudy (anonymous):

so if you can take the derivative, then plug in 13, you will get your answer

OpenStudy (anonymous):

just to finish quick derivative is \[\frac{-16x}{3(1695-8x^2)^{\frac{2}{3}}}\]

OpenStudy (anonymous):

by the chain rule and power rule replace \(x\) by 13 and you should get your answer this is a much snappier way then doing it by hand

OpenStudy (anonymous):

ohh ok thanks!

OpenStudy (anonymous):

yeah i think i got the same

OpenStudy (anonymous):

yw